# An Adaptive Grasping Force Tracking Strategy for Nonlinear and Time-Varying Object Behaviors

### v1

@@ Abstract
Accurate grasp force control is one of the key skills for ensuring successful and damage-free robotic grasping of objects. Although existing methods have conducted in-depth research on slip detection and grasping force planning, they often overlook the issue of adaptive tracking of the actual force to the target force when handling objects with different material properties. The optimal parameters of a force tracking controller are significantly influenced by the object’s stiffness, and many adaptive force tracking algorithms rely on stiffness estimation. However, real-world objects often exhibit viscous, plastic, or other more complex nonlinear time-varying behaviors, and existing studies provide insufficient support for these materials in terms of stiffness definition and estimation. To address this, this paper introduces the concept of generalized stiffness, extending the definition of stiffness to nonlinear time-varying grasp system models, and proposes an online generalized stiffness estimator based on Long Short-Term Memory (LSTM) networks. Based on generalized stiffness, this paper proposes an adaptive parameter adjustment strategy using a PI controller as an example, enabling dynamic force tracking for objects with varying characteristics. Experimental results demonstrate that the proposed method achieves high precision and short probing time, while showing better adaptability to non-ideal objects compared to existing methods. The method effectively solves the problem of grasp force tracking in unknown, nonlinear, and time-varying grasp systems, enhancing the robotic grasping ability in unstructured environments.
@@ I Introduction
Grasping is one of the most important skills for robots. To ensure successful and non-damaging object grasping, it is necessary to precisely control the grasp force applied to the object [1]. In structured environments such as automated production lines, robots are used for repetitive grasping of objects with fixed characteristics, such as weight and shape, and the required grasp force can be pre-calculated. Thus, grasp force control can be achieved through pre-programming or teaching methods [2]. However, with the increasing application of robots in unstructured environments such as daily life and agriculture, where objects vary greatly and are subject to complex environmental constraints (e.g., fruits being pulled by their stems), the unknown object properties and external forces present significant challenges for grasp force control [3, 4].
@@
In terms of execution, force tracking is necessary to ensure that the actual force applied to the object follows the target force dynamically in real time. Since both excessive and insufficient grasp force can damage or cause the object to fall, not only must steady-state errors be kept within a tolerable range, but convergence speed should also be maximized to prevent failure due to overshooting or lag in the actual force. However, many studies [6, 7, 8, 9, 11] focus only on slippage detection or force planning, neglecting force tracking methods. Some studies do mention it, such as Zhou et al. [12] using a PID controller to regulate actual force, Hakkak et al. [13] employing a PI controller, and Muthusam et al. [10] using a fuzzy controller. However, these methods lack adaptability and require offline tuning of optimal controller parameters based on the object’s actual properties to achieve good convergence speed and tracking accuracy. Since object properties are unknown in unstructured environments, controller parameters cannot be predetermined, making it impossible to achieve optimal force tracking performance using these methods. To solve this problem, some research uses sensor data that contains the mechanical characteristics of the object, such as the displacement and force applied to robotic fingers, to estimate system parameters online and adjust the controller accordingly. For example, Treesatayapun [14] proposes a Multi-input Fuzzy Rules Emulated Network (MiFREN) method that can continuously optimize its parameters during grasping. This method can theoretically handle objects with unknown mechanical behaviors, but the network includes 25 parameters that need to be adjusted online, requiring extensive data collection to achieve optimal parameter tuning, leading to longer probing times during initial grasping.
@@
Therefore, to reduce probing time, it is necessary to minimize the number of parameters and focus on the object characteristic parameters that have the most significant impact on the controller. The essence of force tracking is to control the movement of the robotic finger in such a way that the object generates the desired reactive force, and the stiffness of the object reflects the relationship between deformation and applied force. Thus, stiffness is crucial for force tracking controllers. From a biomimetic perspective, this conclusion also holds: humans use multi-sensory coordination to judge the stiffness of an object [15], which has significant implications for human hand grasp control [16]. Many studies adjust the controller by considering the stiffness of the object. De Schutter et al. [17] indicate that object stiffness can effectively characterize system properties and designed a PI controller that adjusts gain based on stiffness, but did not provide a method for online stiffness estimation. Zhang et al. [18] estimate object stiffness based on initial grasp information and adjusts the parameters of a fuzzy logic controller accordingly. Unfortunately, this estimation is only performed during the initial grasp, assuming that the object has fixed stiffness and is a linear, time-invariant system. Once the object stiffness changes (e.g., due to varying grasp forces), the initially estimated stiffness becomes inaccurate, leading to a decline in controller tracking performance. Wei et al. [19] propose a variable impedance force controller that defines object stiffness based on an ideal elastic body model and designs an online stiffness estimator. The impedance parameters of the controller are automatically adjusted according to the estimated stiffness. However, for complex objects with nonlinear or time-varying behaviors, the force-displacement relationship does not follow the ideal elastic body model. If force and displacement data are interpreted according to this model, the stiffness will be incorrectly estimated, causing controller parameter mismatch. In summary, while many studies have considered object stiffness and adjusted controller parameters accordingly, their stiffness definitions and estimation methods are not applicable to more complex nonlinear time-varying systems, resulting in suboptimal performance for real-world objects.
@@ III-A Design of Generalized Stiffness Estimator
Figure 3: Neural network workflow.
The LSTM consists of five layers, each with a hidden dimension of 512. At each time step, the outputs of the two Fully Connected (FC) layers are each one-dimensional. To correct the least squares stiffness estimator, the core idea is to use the output of the LSTM as a correction factor, which is multiplied by the output of the physical heuristic. To eliminate the effect of Batch Normalization (BN) and to avoid the issue where the correction factor processed by the Sigmoid function is less than 1, leading to $\hat{k}(t)$ always being smaller than $\hat{k}_{E}(t)$, we apply a FC layer to $\hat{k}_{E}(t)$ and use ReLU as the activation function to ensure that $\hat{k}(t)>0$ after multiplication.
@@ III-B Data Acquisition, Training and Validation
To train the neural network, we need to acquire input data $[F(t),x(t)]$ and corresponding target outputs $k(t,\mu(t),F(t))$. Since the adaptive grasping scenario in this paper requires the generalized stiffness estimator to work for various objects, we hope the estimator performs well for all systems that meet the assumptions in Chapter II. Thus, we randomly generate 10,000 sets of data that conform to the assumptions in Chapter II (8,000 sets for training and 2,000 for validation). Each set of data simulates a possible grasping process lasting $t=\text{20}\,\text{s}$, with the following mechanisms for generating the various physical quantities:
  • The loading force curve $F(t)$ during the grasping process does not have a fixed form, so it is randomly generated, ensuring it satisfies Assumption 3 from Chapter II. An example of a randomly generated result is shown in Fig. 4(a).
  • The generalized stiffness function $k(t,\mu(t),F)$ is randomly generated. Since both $t$ and $\mu(t)$ affect the stiffness curve in practice, we do not distinguish the individual effects, but ensure the total change is slow and bounded. We generate $k(t,\mu(t),F)=k_{t}(F)$ by first generating a function $k_{0}(F)$, ensuring $0<k_{0}(F)<k_{m}$, and then generating a slowly drifting function $k_{t}(F)$ based on $k_{0}(F)$, while ensuring that the generated result satisfies Assumption 6 from Chapter II.
  • Based on Assumption 4 from Chapter II, we randomly generate a slowly changing zero-force drift function $C(t,\mu(t))=C^{\prime}(t)$.
@@
Given our hardware conditions, we select a control period $T=\text{0.01}\,\text{s}$, with 2,000 time steps. The functions generated earlier are sampled at 0.01s intervals, and we obtain input sequences $[F(t),x(t)]$ and target output sequences $k(t,\mu(t),F(t))=k_{t}(t,F(t))$. By repeating the process described above, multiple sets of simulated grasping data sequences can be obtained. Each randomly generated function has an intuitive physical meaning, and thus can be randomly generated according to the approximate range and distribution characteristics of the various properties of the objects being grasped in real-world scenarios, such as grasp force, zero-force drift, constant-force drift, and stiffness. The distribution of these properties can follow empirical distributions.
@@ V-A Experimental Setup
The force tracking task is completed by a single-degree-of-freedom two-finger robotic gripper, as shown in Fig. 8. The two fingers can move in reverse synchrony along the $Z$-axis. The gripper is equipped with Tac3D tactile sensors and encoders, capable of measuring the force in the $Z$-direction, $F$, the force in the $X-Y$ plane, $F_{T}$, and the position, $x$. The gripper has an STM32 microcontroller, which receives force sensor data, $F$, and encoder readings, $x$, at a frequency of 1kHz and sends speed control commands, $v_{d}$, to the motors. The host computer, a Personal Computer (PC), reads the force and position information from the microcontroller at a frequency of 100Hz and transmits the estimated generalized stiffness $\hat{k}$ and target force $F_{d}$ in real time. The communication diagram is shown in Fig. 8.
@@ V-B Experimental Data Analysis
Figure 10: Experimental results for the paper cup with weight. (a) Actual and target force curves. (b) Stiffness estimation results. (c) Object deformation after force application.
@@
Figure 12: Experimental results for the modeling clay. (a) Actual and target force curves. (b) Stiffness estimation results. (c) Object deformation after force application.
@@ V-C Application Demonstration
Figure 15: Grasping experimental results.
@@
Figure 16: Force tracking results with fixed parameters.
@@
In summary, the generalized stiffness estimation plays a crucial role in our force tracking algorithm and the overall grasping task. Only with accurate generalized stiffness estimates can good force tracking performance be guaranteed, enabling fast grasping and ensuring that the object does not slip or get damaged. This further demonstrates that our force tracking algorithm, which automatically adjusts the parameters based on generalized stiffness, is significantly superior to traditional methods without adaptive parameter tuning.
@@ VI Conclusion
Force tracking experiments on 14 objects show that the proposed method performs well on elastic bodies, viscoelastic bodies, plastic objects, and variable-stiffness objects, with an average asymptotic error of 0.16N, about 1/3 of that achieved by the existing method MiFREN [14], and a probing time of only 3.34s, less than 1/11 of MiFREN’s probing time. Our force tracking algorithm has also shown excellent application results in grasping tasks. For six common non-ideal objects, which are difficult to control in grasping force, successful grasping was achieved through superior force tracking performance. Ablation experiments removing the generalized stiffness estimator demonstrated the irreplaceable role of our stiffness estimator.

### v2

@@ Abstract
Accurate grasp force control is one of the key skills for ensuring successful and damage-free robotic grasping of objects. Although existing methods have conducted in-depth research on slip detection and grasping force planning, they often overlook the issue of adaptive tracking of the actual force to the target force when handling objects with different material properties. The optimal parameters of a force tracking controller are significantly influenced by the object’s stiffness, and many adaptive force tracking algorithms rely on stiffness estimation. However, real-world objects often exhibit viscous, plastic, or other more complex nonlinear time-varying behaviors, and existing studies provide insufficient support for these materials in terms of stiffness definition and estimation. To address this, this paper introduces the concept of generalized stiffness, extending the definition of stiffness to nonlinear time-varying grasp system models, and proposes an online generalized stiffness estimator based on Long Short-Term Memory (LSTM) networks. Based on generalized stiffness, this paper proposes an adaptive parameter adjustment strategy using a PI controller as an example, enabling dynamic force tracking for objects with varying characteristics. Experimental results demonstrate that the proposed method achieves high precision and short probing time, while showing better adaptability to non-ideal objects compared to existing methods. The method effectively solves the problem of grasp force tracking in unknown, nonlinear, and time-varying grasp systems, demonstrating the generalization capability of our neural network and enhancing the robotic grasping ability in unstructured environments.
@@ I Introduction
Grasping is one of the most important skills for robots. To ensure successful and non-damaging object grasping, it is necessary to precisely control the grasp force applied to the object [1]. In structured environments such as automated production lines, robots are used for repetitive grasping of objects with fixed characteristics, such as weight and shape, and the required grasp force can be pre-calculated. Thus, grasp force control can be achieved through pre-programming or teaching methods [2]. Even for challenging soft robotic manipulators, many studies achieve effective control by modeling their dynamics [3, 4, 5], while external forces and object properties remain relatively constant or inapparent, making them easy to model or neglect. However, with the increasing application of robots in unstructured environments such as daily life and agriculture, where objects vary greatly and are subject to complex environmental constraints (e.g., fruits being pulled by their stems), the unknown object properties and external forces present significant challenges for grasp force control [6, 7].
@@
In terms of execution, force tracking is necessary to ensure that the actual force applied to the object follows the target force dynamically in real time. Since both excessive and insufficient grasp force can damage or cause the object to fall, not only must steady-state errors be kept within a tolerable range, but convergence speed should also be maximized to prevent failure due to overshooting or lag in the actual force. However, many studies [9, 10, 11, 12, 14] focus only on slippage detection or force planning, neglecting force tracking methods. Some studies do mention it, such as Zhou et al. [15] using a PID controller to regulate actual force, Hakkak et al. [16] employing a PI controller, and Muthusam et al. [13] using a fuzzy controller. However, these methods lack adaptability and require offline tuning of optimal controller parameters based on the object’s actual properties to achieve good convergence speed and tracking accuracy. Since object properties are unknown in unstructured environments, controller parameters cannot be predetermined, making it impossible to achieve optimal force tracking performance using these methods.
To solve this problem, some research uses sensor data that contains the mechanical characteristics of the object, such as the displacement and force applied to robotic fingers, to estimate system parameters online and adjust the controller accordingly. For example, Treesatayapun [17] proposes a Multi-input Fuzzy Rules Emulated Network (MiFREN) method that can continuously optimize its parameters during grasping. This method can theoretically handle objects with unknown mechanical behaviors, but the network includes 25 parameters that need to be adjusted online, requiring extensive data collection to achieve optimal parameter tuning, leading to longer probing times during initial grasping. Wang et al. [18] propose the Input Convex Long Short-Term Memory Network (ICLSTM), which can predict the next system state given a control attempt, providing valuable information for adaptive control. However, it cannot directly guide the controller’s adjustment. Consequently, a Model Predictive Control (MPC) approach is needed with ICLSTM rollouts at each step, incurring significant overhead and limiting its use to low-frequency applications such as chemical and energy systems.
Therefore, to reduce probing time, it is necessary to minimize the number of parameters and focus on the object characteristic parameters that have the most significant impact on the controller and can guide the controller adjustment in a straightforward manner. Liu et al. [19] and Kumar et al. [20] have employed reinforcement learning to train encoders that automatically extract low-dimensional representations of environmental information for direct use by the control policy; however, in our scenario, such a feature—stiffness—can be explicitly defined. The essence of force tracking is to control the movement of the robotic finger in such a way that the object generates the desired reactive force, and the stiffness of the object reflects the relationship between deformation and applied force. Thus, stiffness is crucial for force tracking controllers. From a biomimetic perspective, this conclusion also holds: humans use multi-sensory coordination to judge the stiffness of an object [21], which has significant implications for human hand grasp control [22].
@@
Many studies adjust the controller by considering the stiffness of the object, but numerous challenges and issues still persist. De Schutter et al. [23] indicate that object stiffness can effectively characterize system properties and designed a PI controller that adjusts gain based on stiffness, but did not provide a method for online stiffness estimation. Zhang et al. [24] estimate object stiffness based on initial grasp information and adjusts the parameters of a fuzzy logic controller accordingly. Unfortunately, this estimation is only performed during the initial grasp, assuming that the object has fixed stiffness and is a linear, time-invariant system. Once the object stiffness changes (e.g., due to varying grasp forces), the initially estimated stiffness becomes inaccurate, leading to a decline in controller tracking performance. Wei et al. [25] propose a variable impedance force controller that defines object stiffness based on an ideal elastic body model and designs an online stiffness estimator. The impedance parameters of the controller are automatically adjusted according to the estimated stiffness. However, for complex objects with nonlinear or time-varying behaviors, the force-displacement relationship does not follow the ideal elastic body model. If force and displacement data are interpreted according to this model, the stiffness will be incorrectly estimated, causing controller parameter mismatch. In summary, while many studies have considered object stiffness and adjusted controller parameters accordingly, their stiffness definitions and estimation methods are not applicable to more complex nonlinear time-varying systems, resulting in suboptimal performance for real-world objects. In this context, Wu et al. [26] inspired us by demonstrating the powerful capability of neural networks to estimate specific object parameters from observations of nonlinear systems. However, unlike the object mass estimated in this work, the stiffness of many objects is highly time-varying, which necessitates a more sophisticated design to enable neural networks to perform real-time and accurate estimation.
@@ III-A Design of Generalized Stiffness Estimator
Figure 3: Neural network workflow. $F$ is the grasping force, $x$ is the finger displacement, $\hat{k}_{E}$ is the stiffness estimated by LSM, and $\hat{k}$ is the generalized stiffness estimated by our method.
The LSTM consists of five layers, each with a hidden dimension of 512. Additionally, the network includes two Fully Connected (FC) layers. The first FC layer has an input dimension of 512 and, at each time step, receives the hidden state from the final LSTM cell, producing a one-dimensional vector as its output. The second FC layer, with both input and output dimensions of 1, serves to scale the output of the LSM. To correct the least squares stiffness estimator, the core idea is to use the output of the LSTM (and the FC layer behind) as a correction factor, which is multiplied by the output of the physical heuristic. To eliminate the effect of Batch Normalization (BN) and to avoid the issue where the correction factor processed by the Sigmoid function is less than 1, leading to $\hat{k}(t)$ always being smaller than $\hat{k}_{E}(t)$, we apply a FC layer to $\hat{k}_{E}(t)$ and use ReLU as the activation function to ensure that $\hat{k}(t)>0$ after multiplication.
@@ III-B Data Acquisition, Training and Validation
To train the neural network, we need to acquire input data $[F(t),x(t)]$ and corresponding target outputs $k(t,\mu(t),F(t))$. Since the adaptive grasping scenario in this paper requires the generalized stiffness estimator to work for various objects, we hope the estimator performs well for all systems that meet the assumptions in Chapter II, where we defined a comprehensive system model that nearly encompasses the wide array of nonlinear time-varying grasping processes observed in practice, using Equation (4) along with specific constraints on the physical quantities. Consequently, almost all real-world grasping force tracking processes—including variations in force, displacement, and the object’s generalized stiffness—fall within this model.
Therefore, we generate the training data by randomly sampling multiple grasping processes from this system model using an empirical distribution. Any real-world force-tracking process that adheres to our system model will benefit from interpolation generalization—even if its distribution deviates slightly from the empirical one. This interpolation generalization is relatively straightforward for neural networks, and is also tested in Chapter V.
In our experiment, the training data were generated by randomly sampling 8,000 grasping processes from this system model based on an empirical distribution, and an additional 2,000 sets (drawn independently but sharing the same distribution) were used for validation. Each dataset simulates a possible grasping process lasting $t=\text{20}\,\text{s}$, with the following mechanisms for generating the various physical quantities:
  • The loading force curve $F(t)$ during the grasping process does not have a fixed form, so it is randomly generated, ensuring it satisfies Assumption 3 from Chapter II. Empirically, the grasping force appears to be approximately uniformly distributed in the range of 0.1N to 10N, so we adopt a uniform distribution as an empirical approximation for $F$. An example of a randomly generated result is shown in Fig. 4(a).
  • The generalized stiffness function $k(t,\mu(t),F)$ is randomly generated. Since both $t$ and $\mu(t)$ affect the stiffness curve in practice, we do not distinguish their individual effects, but ensure the total change is slow and bounded. We generate $k(t,\mu(t),F)=k_{t}(F)$ by first generating a function $k_{0}(F)$, ensuring $0<k_{0}(F)<k_{m}$, and then generating a slowly drifting function $k_{t}(F)$ based on $k_{0}(F)$, while ensuring that the generated result satisfies Assumption 6.
    There are multiple ways to generate the bivariate function $k_{t}(F)$, but it is important to note that a simple uniform distribution does not align with empirical observations, as it would result in a low probability of selecting objects with small stiffness. In our scenario, considering Assumption 2 and real-world objects, the stiffness has an upper bound of 8N/mm and a lower bound of 0.05N/mm. Therefore, we model the logarithm of stiffness as uniformly distributed, i.e., $\ln k_{t}(F)\sim\mathcal{U}(\ln(0.05),\ln(8))$, which better reflects the real-world distribution of various object stiffness levels.
[…]
@@
Given our hardware conditions, we select a control period $T=\text{0.01}\,\text{s}$, with 2,000 time steps. The functions generated earlier are sampled at 0.01s intervals, and we obtain input sequences $[F(t),x(t)]$ and target output sequences $k(t,\mu(t),F(t))=k_{t}(F(t))$. By repeating the process described above, multiple sets of simulated grasping data sequences can be obtained. Each randomly generated function has an intuitive physical meaning, and thus can be randomly generated according to the approximate range and distribution characteristics of the various properties of the objects being grasped in real-world scenarios, such as grasp force, zero-force drift, constant-force drift, and stiffness. The distribution of these properties can follow empirical distributions, and random functions can be obtained in various ways, such as selecting characteristic points and performing interpolation.
@@ V-A Experimental Setup
The force tracking task is completed by a single-degree-of-freedom two-finger robotic gripper, as shown in Fig. 8. The two fingers can move in reverse synchrony along the $Z$-axis. The gripper is equipped with Tac3D tactile sensors and encoders, capable of measuring the force in the $Z$-direction, $F$, the force in the $X-Y$ plane, $F_{T}$, and the position, $x$. In fact, although this article focuses on the grasping force tracking, numerous studies have demonstrated that vision-based tactile sensors are capable of measuring more complex force information and guiding the setting of target grasping forces [27, 28]. By combining these capabilities, a comprehensive grasping force control system can be achieved. The gripper has an STM32 microcontroller, which receives force sensor data, $F$, and encoder readings, $x$, at a frequency of 1kHz and sends speed control commands, $v_{d}$, to the motors. The host computer, a Personal Computer (PC) powered by an AMD Ryzen 7 6800H CPU without GPU acceleration, reads the force and position information from the microcontroller at a frequency of 100Hz, performs real-time stiffness estimation and control computations, and transmits the estimated generalized stiffness $\hat{k}$ and target force $F_{d}$ in real time. The communication diagram is shown in Fig. 8.
@@ V-B Experimental Data Analysis
Figure 10: Experimental results for the paper cup with weight. (a) Curves of the actual force and the target force $F_{d}$. (b) Stiffness estimation results. (c) Object deformation after force application.
@@
Figure 12: Experimental results for the modeling clay. (a) Curves of the actual force and the target force $F_{d}$. (b) Stiffness estimation results. (c) Object deformation after force application.
@@
To further understand the performance of our proposed method, we analyze the primary sources of error in our system:
  • Limited performance of mechanical actuators: Our force control is based on the built-in velocity loop of the robotic manipulator’s motors, which introduces velocity control errors. As a result, the manipulator cannot perfectly follow our motion commands. More precise actuators would help reduce the existing errors.
  • Theoretical limitations of the control model: Since our force control relies on a PI controller, as shown in Figure 7, the system matrix has a minimum spectral norm of 0.707. This implies that the system still requires several time steps to converge the error close to zero. Adopting a control strategy with faster convergence than PI control as the underlying framework would help accelerate the error convergence.
  • Errors in stiffness estimation: Stiffness estimation is performed at a frequency of 100 Hz and is based on predicting future stiffness changes using historical information. Abrupt changes in the object’s stiffness (e.g., a paper cup with weight) can cause the estimated stiffness to deviate from its true value, leading to a deviation of the system matrix’s spectral norm from its minimum and slowing down the convergence speed. Enhancing stiffness estimation algorithms may help speed up error convergence.
@@ V-C Application Demonstration
Figure 15: Grasping experimental results. $F$ is the grasping force, $F_{d}$ is the target force, and $F_{T}$ is the tangential force.
@@
Figure 16: Force tracking results with fixed parameters. $F$ is the grasping force, $F_{d}$ is the target force, and $F_{T}$ is the tangential force.
@@
These results are consistent with the theoretical analysis derived from Equation (21), which indicates that an excessively large $\hat{k}$ results in smaller effective PI coefficients, leading to slow convergence. Conversely, a small $\hat{k}$ can cause the PI parameters to become too large, resulting in overshooting or oscillation. In summary, the generalized stiffness estimation plays a crucial role in our force tracking algorithm and the overall grasping task. Only with accurate generalized stiffness estimates can good force tracking performance be guaranteed, enabling fast grasping and ensuring that the object does not slip or get damaged. This further demonstrates that our force tracking algorithm, which automatically adjusts the parameters based on generalized stiffness, is significantly superior to traditional methods without adaptive parameter tuning.
@@ VI Conclusion
Force tracking experiments on 14 objects show that the proposed method performs well on elastic bodies, viscoelastic bodies, plastic objects, and variable-stiffness objects, with an average asymptotic error of 0.16N, about 1/3 of that achieved by the existing method MiFREN [17], and a probing time of only 3.34s, less than 1/11 of MiFREN’s probing time. Our force tracking algorithm has also shown excellent application results in grasping tasks. For six common non-ideal objects, which are difficult to control in grasping force, successful grasping was achieved through superior force tracking performance. These experiments demonstrate successful generalization of our neural network from training data to real-world applications. Ablation experiments removing the generalized stiffness estimator demonstrated the irreplaceable role of our stiffness estimator.